# HYDRA: Unifying Multi-modal Generation and Understanding via Representation-Harmonized Tokenization

### 16 Mar 2026 v1

@@ 3 Experiments
We evaluate HYDRA at two scales: 1.5B (based on Qwen2.5-1.5B-Instruct (qwen2.5)) and 7B (Qwen2.5-7B-Instruct). The HYDRA-TOK is initialized with InternViT-2.5 (internvl2.5) to leverage robust visual priors. The training recipe employs a three-stage progressive pipeline optimized with AdamW (loshchilov2017decoupled).The specific learning rate schedules for each stage are detailed in Tab. 6. For comprehensive experimental details and evaluation metrics regarding all ablation studies, please refer to Appendix D.1.
@@ 3.1 Main Results
We next verify whether our pure ViT-based HYDRA-TOK can serve as a high-quality visual foundation. Historically, ViTs were deemed unsuitable for dense pixel-level synthesis compared to Convolutional VAEs. Tab. 1 challenges this assumption. HYDRA-TOK achieves a PSNR of 36.39 and an rFID of 0.08 on ImageNet-1K (imagenet). This performance significantly outperforms unified baselines like BLIP3-o (14.71 PSNR) (blip3-o) and even surpasses specialized generative tokenizers such as FLUX-VAE (32.74 PSNR) (flux). This confirms that our functionally progressive design (specifically the Gen-ViT) successfully captures compact structure-preserving primitives, allowing a pure ViT to supersede traditional VAEs as a universal visual tokenizer.
@@
We assess the coherence of information flow by evaluating multimodal understanding performance, as shown in Tab. 2. Sequential architectures, such as Show-o2 (xie2025show), often compromise information coherence due to the compression bottleneck imposed by VAEs. In contrast, HYDRA maintains a continuous and uncompressed information flow throughout its pure ViT backbone. At the 1.5B scale, HYDRA demonstrates dominant performance with an average score of 63.1, surpassing the strong baseline Show-o2 (53.2) by a significant margin. This advantage is particularly pronounced in fine-grained tasks sensitive to information loss, such as OCRBench (liu2024ocrbench), where HYDRA achieves 50.8, a score more than double that of Show-o2 (24.5). Scaling up to 7B, HYDRA continues to lead, outperforming Show-o2 on complex reasoning benchmarks like MMStar (chen2024mmstar) (62.3 vs. 56.6) and SEED (li2023seed) (75.5 vs. 69.8). Most notably, in OCRBench (liu2024ocrbench), HYDRA preserves high-frequency character details often discarded by VAE-based unified models, scoring 57.7 compared to 32.4 for Show-o2.
@@
Supported by this harmonized training dynamic, we further validate the learning compatibility on text-to-image generation benchmarks in Tab. 3. A common failure mode in UMMs is the “tug-of-war” where improving understanding degrades generation. HYDRA breaks this trade-off, demonstrating that joint optimization leads to superior generative performance. At the 1.5B scale, HYDRA establishes a new benchmark for native UMMs, achieving an Overall GenEval score of 0.86 and DPG-Bench score of 85.51, significantly outperforming Show-o2 (0.73 / 85.02). At the 7B scale, HYDRA sets new state-of-the-art records with a GenEval Overall score of 0.86, surpassing both the unified model Ming-UniVision (huang2025ming) (0.85) and the specialized 12B model FLUX.1 [Dev] (0.82) (flux). Furthermore, on the WISE benchmark, HYDRA achieves an overall score of 0.53, demonstrating robust alignment across diverse cultural and spatial contexts compared to existing native unified baselines.
@@ A.1 Unified Multimodal Models
The pursuit of a singular architecture for both visual understanding and generation has led to the evolution of Unified Multimodal Models (UMMs) (liao2025mogao; deng2025emerging; hu2025omni; li2025unif), broadly categorized into composite and native architectures.
Composite UMMs such as MetaQuery (metamorph), BLIP-3o (blip3-o), OpenUni (wu2025openuni) and Unilip (unilip) achieve capability extension by integrating specialized, pre-trained experts (e.g., separate LLMs and Diffusion models) via lightweight adapters. While effective for rapid deployment, their reliance on frozen backbones restricts deep inter-modal interaction, resulting in a shallow unification where understanding and generation remain functionally decoupled
Native UMMs strive for deeper integration by training holistically within a shared parameter space. However, as analyzed in Fig. 1, existing methods are constrained by a fundamental representation trilemma: (i) Decoupled Architectures, such as the Janus series (wu2025janus; ma2025janusflow), Bagel (bagel) and UniToken (jiao2025unitoken), mitigate task interference by utilizing separate or dual encoders for vision and language. While UniToken concatenates discrete and continuous tokens to bridge modalities, this dual-encoder design inherently sacrifices Unification of Input Representation, leading to parameter redundancy, computational inefficiency due to sequence inflation, and severed task synergy. (ii) Sequential Architectures, including Show-o2 (xie2025show) and TUNA (liu2025tuna), attempt to unify inputs by stacking semantic encoders atop a VAE. However, the heavy compression required for the VAE acts as an information bottleneck, disrupting the Coherence of Information Flow and diluting the fine-grained structural signals needed for generation. (iii) Shared Architectures, such as Unitok (unitok) and Transfusion (zhou2024transfusion; wu2025harmonizing), attempt to unify tasks but struggle with the Compatibility of Learning Process. On one hand, discrete approaches like Unitok (unitok) suffer from quantization errors inherent in codebooks, which fundamentally conflict with the high-fidelity requirements of synthesis. On the other hand, continuous approaches like Transfusion (zhou2024transfusion) and Harmon (wu2025harmonizing) employ VAE latent features for both tasks; while these representations are adequate for generation, they lack the high-dimensional semantic abstractions required for perception, resulting in suboptimal performance on complex understanding tasks. In contrast, our HYDRA achieves rational unification by resolving these conflicts through a harmonized, continuous representation.
@@ A.2 Unified Tokenizers and Representations
Recent research has shifted from discrete to continuous unified representations to bridge the gap between semantic abstraction and pixel-level fidelity. To circumvent the traditional variational bottleneck, paradigms such as RAE (rae) and SVG (shi2025latent) employ frozen, pre-trained semantic encoders. While this accelerates convergence, the inherent loss of low-level features compromises reconstruction fidelity. Although UniFlow (yue2025uniflow) attempts to recover structural details via self-distillation, it relies on dual-stream features, failing to achieve a truly unified, single-stream information flow. Similarly, QLIP (qlip) seeks to unify representations by aligning visual tokens with textual semantics; however, this heavy reliance on linguistic alignment imposes a semantic bottleneck, filtering out the high-frequency textural details necessary for photorealistic synthesis. Furthermore, while VTP (yao2025vtp) improves generative scaling by jointly optimizing contrastive and reconstruction losses, its low-dimensional latent space is tailored for generation and lacks the representational capacity required for complex multimodal perception. Distinctly, our HYDRA-TOK employs a functionally progressive pure ViT backbone to actively bridge the representation gap. This design ensures a coherent single-stream flow, eliminating quantization errors while simultaneously satisfying the conflicting demands of high-fidelity synthesis and deep semantic understanding.
@@ C.2 Scaling HYDRA-TOK training data
For multimodal understanding capabilities, the tokenizer is trained utilizing the LLaVA-1.5 setting (llava1.5). As illustrated in Fig. 6(a), we evaluate performance across three data regimes: 1.2M, 4M, and 20M image-text pairs. The evaluation metrics include generation (Geneval), reconstruction (rFID), and understanding (Avg QA), where Avg QA denotes the average score across the POPE (pope), MMBench (mmbench), MMMU (yue2024mmmu), AI2D (kembhavi2016ai2d), and RealWorldQA benchmarks. We observe distinct trends for different capabilities as the data size increases. The generative capability, indicated by the Geneval score, shows a consistent positive trend, improving steadily from approximately 38 at 1.2M to over 45 when trained on 20M data. Reconstruction fidelity, measured by rFID (where lower is better), also benefits significantly from larger data scales. While it shows a slight increase at 4M, it achieves its best performance with a sharp drop to approximately 0.08 at the 20M mark. Conversely, multimodal understanding capabilities, as reflected by the Avg QA score, remain relatively stable and robust across all data sizes, maintaining a score consistently above 61. These results suggest that while understanding capabilities are established early, scaling the tokenizer training data is critical for optimizing generation and reconstruction performance.
@@
Figure 7: Analysis of tokenizer data scaling and layer-wise representations. (a) This figure illustrates how understanding (Avg QA), generation (Geneval), and reconstruction (rFID) metrics evolve as the tokenizer training data size increases from 1.2M to 20M. (b) The figure shows the CKNNA score across the 24 layers of the HYDRA-TOK, indicating the change in representational similarity between teacher vit (internvl2.5).
@@ C.3 Visualization of CKNNA
To better illustrate the coherence of information flow within our model, we conducted a representational similarity analysis using Centered Kernel Nearest-Neighbor Alignment (CKNNA) (huh2024platonic). We randomly selected 10,000 images from the ImageNet2012 validation set (imagenet) to calculate the CKNNA metric between our HYDRA-TOK and the teacher model, InternViT. As observed in Fig. 6(b), our model exhibits strong alignment in early layers.
@@ C.4 Visualization of t-SNE
As shown in Fig. 8, we visualize the learned features from both the generation and understanding branches of HYDRA, comparing them against UniFlow (yue2025uniflow) and Show-o2 (xie2025show) (equipped with WAN (wan2.2) and SigLIP (tschannen2025siglip)). While the baselines often exhibit disparately distributed features, HYDRA demonstrates distinct class clusters in both its generation and understanding representations. This strong semantic discriminability indicates a high degree of alignment between the two feature spaces. Such similarity confirms that our architecture establishes a coherent information flow, enabling the two tasks to be collaboratively optimized within a harmonized representational framework.
@@ D.1 Ablation study training details
In our ablation studies, the evaluation covers three core capabilities with specific setups: (i) Multimodal Understanding: Following the LLaVA-1.5 training protocol (llava1.5), we report the average score (Avg QA) across POPE (pope), MMBench (mmbench), MMMU (yue2024mmmu), AI2D (kembhavi2016ai2d), and RealWorldQA. (ii) Image Generation: We adapt the Nitro-E framework (nitro) by replacing its VAE with our HYDRA-TOK. We train this setup for 15k iterations using the SAM-1B (sam) and JourneyDB (sun2023journeydb) datasets and evaluate performance on GenEval. (iii) Image Reconstruction: We train on the ImageNet-1k (1.2M) dataset (imagenet) for 150k iterations and assess quality using rFID.
@@ D.3.1 Image Understanding
For the initial pre-training stage, we utilized a massive dataset comprising approximately 22M images paired with captions. These data were sourced from a diverse collection of large-scale datasets, including ImageNet (imagenet), DataComp-1B (gadre2023datacomp), COYO-700M (kakaobrain2022coyo-700m), SA-1B (sam), and DenseFusion (li2024densefusion).
@@
In Stage II, we continued to utilize the comprehensive collection of data from the sources employed in Stage I (ImageNet (imagenet), DataComp-1B (gadre2023datacomp), COYO-700M (kakaobrain2022coyo-700m), SA-1B (sam), and DenseFusion (li2024densefusion)) for further training.
@@
For the final high-quality fine-tuning stage, we directly utilized the refined 3.2M instruction-tuning data from LLaVA-One-Vision (Llava-onevision) to maximize image understanding performance.
@@ D.3.2 Image Generation
For the initial generation training stage, we curated a large-scale dataset totaling approximately 78M samples. This comprises SA-1B (sam), CC-12M (changpinyo2021conceptual), 1M images from DenseFusion (li2024densefusion), and a internal dataset consisting of 55M real images.
@@
Stage II expands upon the data used in Stage I. In addition to SA-1B (sam), CC-12M (changpinyo2021conceptual), the 55M internal real images, and DenseFusion-1M (li2024densefusion), we incorporated 10M synthetic images generated through Flux distillation, following the methodology described in (nitro).
@@
For high-quality generation fine-tuning, we utilized the JourneyDB dataset (sun2023journeydb) and DenseFusion-1M (li2024densefusion). Additionally, we refined the 10M Flux-distilled data from Stage II through filtering, selecting a high-quality subset of 5M images for this final stage.
@@ Appendix E Additional main results
Table 7: Detailed image generation results on the GenEval benchmark. refers to fintuning with GPT-4o distilled dataset (blip3-o).
@@
Models Size Single Object Two Objects Count Colors Position Color Attribute Overall Generation-only Models SD3-Med (esser2024scaling) 2B 0.99 0.94 0.72 0.89 0.33 0.60 0.74 FLUX.1 [Dev] (flux) 12B 0.98 0.93 0.75 0.93 0.68 0.65 0.82 DALL-E 3 (dalle) - 0.96 0.87 0.47 0.83 0.43 0.45 0.67 Composite Unified Multimodal Models TokenFlow-XL (qu2025tokenflow) 14B 0.95 0.60 0.41 0.81 0.16 0.24 0.55 SEED-X (seed-x) 17B 0.97 0.58 0.26 0.80 0.19 0.14 0.49 MetaQuery-XL (pan2025transfer) 7B - - - - - - 0.80 Blip3-o (blip3-o) 8B - - - - - - 0.84 Smaller Native Unified Multimodal Models Show-o (xie2024show) 1.3B 0.95 0.52 0.49 0.82 0.11 0.28 0.53 Harmon (wu2025harmonizing) 1.5B 0.99 0.86 0.66 0.85 0.74 0.48 0.76 Show-o2 (xie2025show) 1.5B 0.99 0.86 0.55 0.86 0.46 0.63 0.73 HYDRA 1.5B 1.00 0.95 0.75 0.91 0.76 0.77 0.86 Larger Native Unified Multimodal Models Ming-UniVision (huang2025ming) 16B 1.00 0.93 0.59 0.93 0.92 0.70 0.85 BAGEL (deng2025emerging) 14B 0.98 0.95 0.84 0.95 0.78 0.77 0.88 MUSE-VL (muse-vl) 7B 0.98 0.64 0.54 0.72 0.25 0.31 0.57 Janus-Pro (chen2025janus) 7B 0.99 0.89 0.59 0.90 0.79 0.66 0.80 Show-o2 (xie2025show) 7B 1.00 0.87 0.58 0.92 0.52 0.62 0.76 HYDRA 7B 1.00 0.97 0.68 0.91 0.81 0.80 0.86
Table 8: Detailed image generation results on the DPG-Bench benchmark (hu2024ella).
@@
Models Size Global Entity Attribute Relation Other Overall Generation-only Models FLUX.1 [Dev] (flux) 12B 82.10 89.50 88.70 91.10 89.40 84.00 Smaller Native Unified Multimodal Models Show-o2 (xie2025show) 1.5B 87.53 90.38 91.34 90.30 91.21 85.02 HYDRA 1.5B 89.65 91.62 90.24 91.59 90.88 85.51 Larger Native Unified Multimodal Models Janus-Pro (chen2025janus) 7B 86.90 88.90 89.40 89.32 89.48 84.19 BAGEL (deng2025emerging) 14B 88.94 90.37 91.29 90.82 88.67 85.07 Show-o2 (xie2025show) 7B 89.00 91.78 89.96 91.81 91.64 86.14 HYDRA 7B 90.22 91.88 89.92 91.83 90.93 86.40
Table 9: Detailed image generation results on the WISE benchmark (niu2025wise).
@@
Models Size Culture Time Space Biology Physics Chemistry Overall Generation-only Models FLUX.1 [Dev] (flux) 12B 0.48 0.58 0.62 0.42 0.51 0.35 0.50 SD3.5-Large (esser2024scaling) 8B 0.44 0.50 0.58 0.44 0.52 0.31 0.46 Composite Unified Multimodal Models Blip3-o-4B (blip3-o) 4B - - - - - - 0.50 Smaller Native Unified Multimodal Models Harmon (wu2025harmonizing) 1.5B 0.38 0.48 0.52 0.37 0.44 0.29 0.41 Show-o2 (xie2025show) 1.5B 0.33 0.41 0.53 0.35 0.49 0.31 0.39 HYDRA 1.5B 0.49 0.47 0.64 0.42 0.59 0.34 0.49 Larger Native Unified Multimodal Models BAGEL (deng2025emerging) 14B 0.44 0.55 0.68 0.44 0.60 0.39 0.52 VILA-U-7B (wu2024vila) 7B 0.26 0.33 0.37 0.35 0.39 0.23 0.31 Janus-Pro-7B (chen2025janus) 7B 0.30 0.37 0.49 0.36 0.42 0.26 0.35 Emu3-Gen-8B (wang2024emu3) 8B 0.34 0.45 0.48 0.41 0.45 0.27 0.39 Show-o2 (xie2025show) 7B 0.40 0.45 0.58 0.39 0.53 0.34 0.44 HYDRA 7B 0.52 0.53 0.68 0.47 0.58 0.38 0.53

### 17 Mar 2026 v2

@@ 3 Experiments
We evaluate HYDRA at two scales: 1.5B (based on Qwen2.5-1.5B-Instruct (Yang et al., 2024)) and 7B (Qwen2.5-7B-Instruct). The HYDRA-TOK is initialized with InternViT-2.5 (Chen et al., 2024b) to leverage robust visual priors. The training recipe employs a three-stage progressive pipeline optimized with AdamW (Loshchilov & Hutter, 2017).The specific learning rate schedules for each stage are detailed in Tab. 6. For comprehensive experimental details and evaluation metrics regarding all ablation studies, please refer to Appendix D.1.
@@ 3.1 Main Results
We next verify whether our pure ViT-based HYDRA-TOK can serve as a high-quality visual foundation. Historically, ViTs were deemed unsuitable for dense pixel-level synthesis compared to Convolutional VAEs. Tab. 1 challenges this assumption. HYDRA-TOK achieves a PSNR of 36.39 and an rFID of 0.08 on ImageNet-1K (Russakovsky et al., 2014). This performance significantly outperforms unified baselines like BLIP3-o (14.71 PSNR) (Chen et al., 2025a) and even surpasses specialized generative tokenizers such as FLUX-VAE (32.74 PSNR) (Labs et al., 2025). This confirms that our functionally progressive design (specifically the Gen-ViT) successfully captures compact structure-preserving primitives, allowing a pure ViT to supersede traditional VAEs as a universal visual tokenizer.
@@
We assess the coherence of information flow by evaluating multimodal understanding performance, as shown in Tab. 2. Sequential architectures, such as Show-o2 (Xie et al., 2025a), often compromise information coherence due to the compression bottleneck imposed by VAEs. In contrast, HYDRA maintains a continuous and uncompressed information flow throughout its pure ViT backbone. At the 1.5B scale, HYDRA demonstrates dominant performance with an average score of 63.1, surpassing the strong baseline Show-o2 (53.2) by a significant margin. This advantage is particularly pronounced in fine-grained tasks sensitive to information loss, such as OCRBench (Liu et al., 2024b), where HYDRA achieves 50.8, a score more than double that of Show-o2 (24.5). Scaling up to 7B, HYDRA continues to lead, outperforming Show-o2 on complex reasoning benchmarks like MMStar (Chen et al., 2024a) (62.3 vs. 56.6) and SEED (Li et al., 2023a) (75.5 vs. 69.8). Most notably, in OCRBench (Liu et al., 2024b), HYDRA preserves high-frequency character details often discarded by VAE-based unified models, scoring 57.7 compared to 32.4 for Show-o2.
@@
Supported by this harmonized training dynamic, we further validate the learning compatibility on text-to-image generation benchmarks in Tab. 3. A common failure mode in UMMs is the “tug-of-war” where improving understanding degrades generation. HYDRA breaks this trade-off, demonstrating that joint optimization leads to superior generative performance. At the 1.5B scale, HYDRA establishes a new benchmark for native UMMs, achieving an Overall GenEval score of 0.86 and DPG-Bench score of 85.51, significantly outperforming Show-o2 (0.73 / 85.02). At the 7B scale, HYDRA sets new state-of-the-art records with a GenEval Overall score of 0.86, surpassing both the unified model Ming-UniVision (Huang et al., 2025) (0.85) and the specialized 12B model FLUX.1 [Dev] (0.82) (Labs et al., 2025). Furthermore, on the WISE benchmark, HYDRA achieves an overall score of 0.53, demonstrating robust alignment across diverse cultural and spatial contexts compared to existing native unified baselines.
@@ A.1 Unified Multimodal Models
The pursuit of a singular architecture for both visual understanding and generation has led to the evolution of Unified Multimodal Models (UMMs) (Liao et al., 2025a; Deng et al., 2025b; Hu et al., 2025; Li et al., 2025), broadly categorized into composite and native architectures.
Composite UMMs such as MetaQuery (Gupta et al., 2022), BLIP-3o (Chen et al., 2025a), OpenUni (Wu et al., 2025c) and Unilip (Tang et al., 2025) achieve capability extension by integrating specialized, pre-trained experts (e.g., separate LLMs and Diffusion models) via lightweight adapters. While effective for rapid deployment, their reliance on frozen backbones restricts deep inter-modal interaction, resulting in a shallow unification where understanding and generation remain functionally decoupled
Native UMMs strive for deeper integration by training holistically within a shared parameter space. However, as analyzed in Fig. 1, existing methods are constrained by a fundamental representation trilemma: (i) Decoupled Architectures, such as the Janus series (Wu et al., 2025b; Ma et al., 2025c), Bagel (Deng et al., 2025a) and UniToken (Jiao et al., 2025), mitigate task interference by utilizing separate or dual encoders for vision and language. While UniToken concatenates discrete and continuous tokens to bridge modalities, this dual-encoder design inherently sacrifices Unification of Input Representation, leading to parameter redundancy, computational inefficiency due to sequence inflation, and severed task synergy. (ii) Sequential Architectures, including Show-o2 (Xie et al., 2025a) and TUNA (Liu et al., 2025), attempt to unify inputs by stacking semantic encoders atop a VAE. However, the heavy compression required for the VAE acts as an information bottleneck, disrupting the Coherence of Information Flow and diluting the fine-grained structural signals needed for generation. (iii) Shared Architectures, such as Unitok (Ma et al., 2025b) and Transfusion (Zhou et al., 2024; Wu et al., 2025d), attempt to unify tasks but struggle with the Compatibility of Learning Process. On one hand, discrete approaches like Unitok (Ma et al., 2025b) suffer from quantization errors inherent in codebooks, which fundamentally conflict with the high-fidelity requirements of synthesis. On the other hand, continuous approaches like Transfusion (Zhou et al., 2024) and Harmon (Wu et al., 2025d) employ VAE latent features for both tasks; while these representations are adequate for generation, they lack the high-dimensional semantic abstractions required for perception, resulting in suboptimal performance on complex understanding tasks. In contrast, our HYDRA achieves rational unification by resolving these conflicts through a harmonized, continuous representation.
@@ A.2 Unified Tokenizers and Representations
Recent research has shifted from discrete to continuous unified representations to bridge the gap between semantic abstraction and pixel-level fidelity. To circumvent the traditional variational bottleneck, paradigms such as RAE (Zheng et al., 2025) and SVG (Shi et al., 2025) employ frozen, pre-trained semantic encoders. While this accelerates convergence, the inherent loss of low-level features compromises reconstruction fidelity. Although UniFlow (Yue et al., 2025) attempts to recover structural details via self-distillation, it relies on dual-stream features, failing to achieve a truly unified, single-stream information flow. Similarly, QLIP (Zhao et al., 2025) seeks to unify representations by aligning visual tokens with textual semantics; however, this heavy reliance on linguistic alignment imposes a semantic bottleneck, filtering out the high-frequency textural details necessary for photorealistic synthesis. Furthermore, while VTP (Yao et al., 2025a) improves generative scaling by jointly optimizing contrastive and reconstruction losses, its low-dimensional latent space is tailored for generation and lacks the representational capacity required for complex multimodal perception. Distinctly, our HYDRA-TOK employs a functionally progressive pure ViT backbone to actively bridge the representation gap. This design ensures a coherent single-stream flow, eliminating quantization errors while simultaneously satisfying the conflicting demands of high-fidelity synthesis and deep semantic understanding.
@@ C.2 Scaling HYDRA-TOK training data
For multimodal understanding capabilities, the tokenizer is trained utilizing the LLaVA-1.5 setting (Liu et al., 2023a). As illustrated in Fig. 6(a), we evaluate performance across three data regimes: 1.2M, 4M, and 20M image-text pairs. The evaluation metrics include generation (Geneval), reconstruction (rFID), and understanding (Avg QA), where Avg QA denotes the average score across the POPE (Li et al., 2023b), MMBench (Liu et al., 2024a), MMMU (Yue et al., 2024), AI2D (Kembhavi et al., 2016), and RealWorldQA benchmarks. We observe distinct trends for different capabilities as the data size increases. The generative capability, indicated by the Geneval score, shows a consistent positive trend, improving steadily from approximately 38 at 1.2M to over 45 when trained on 20M data. Reconstruction fidelity, measured by rFID (where lower is better), also benefits significantly from larger data scales. While it shows a slight increase at 4M, it achieves its best performance with a sharp drop to approximately 0.08 at the 20M mark. Conversely, multimodal understanding capabilities, as reflected by the Avg QA score, remain relatively stable and robust across all data sizes, maintaining a score consistently above 61. These results suggest that while understanding capabilities are established early, scaling the tokenizer training data is critical for optimizing generation and reconstruction performance.
@@
Figure 7: Analysis of tokenizer data scaling and layer-wise representations. (a) This figure illustrates how understanding (Avg QA), generation (Geneval), and reconstruction (rFID) metrics evolve as the tokenizer training data size increases from 1.2M to 20M. (b) The figure shows the CKNNA score across the 24 layers of the HYDRA-TOK, indicating the change in representational similarity between teacher vit (Chen et al., 2024b).
@@ C.3 Visualization of CKNNA
To better illustrate the coherence of information flow within our model, we conducted a representational similarity analysis using Centered Kernel Nearest-Neighbor Alignment (CKNNA) (Huh et al., 2024). We randomly selected 10,000 images from the ImageNet2012 validation set (Russakovsky et al., 2014) to calculate the CKNNA metric between our HYDRA-TOK and the teacher model, InternViT. As observed in Fig. 6(b), our model exhibits strong alignment in early layers.
@@ C.4 Visualization of t-SNE
As shown in Fig. 8, we visualize the learned features from both the generation and understanding branches of HYDRA, comparing them against UniFlow (Yue et al., 2025) and Show-o2 (Xie et al., 2025a) (equipped with WAN (Wan et al., 2025) and SigLIP (Tschannen et al., 2025)). While the baselines often exhibit disparately distributed features, HYDRA demonstrates distinct class clusters in both its generation and understanding representations. This strong semantic discriminability indicates a high degree of alignment between the two feature spaces. Such similarity confirms that our architecture establishes a coherent information flow, enabling the two tasks to be collaboratively optimized within a harmonized representational framework.
@@ D.1 Ablation study training details
In our ablation studies, the evaluation covers three core capabilities with specific setups: (i) Multimodal Understanding: Following the LLaVA-1.5 training protocol (Liu et al., 2023a), we report the average score (Avg QA) across POPE (Li et al., 2023b), MMBench (Liu et al., 2024a), MMMU (Yue et al., 2024), AI2D (Kembhavi et al., 2016), and RealWorldQA. (ii) Image Generation: We adapt the Nitro-E framework (Shen et al., 2025) by replacing its VAE with our HYDRA-TOK. We train this setup for 15k iterations using the SAM-1B (Kirillov et al., 2023) and JourneyDB (Sun et al., 2023) datasets and evaluate performance on GenEval. (iii) Image Reconstruction: We train on the ImageNet-1k (1.2M) dataset (Russakovsky et al., 2014) for 150k iterations and assess quality using rFID.
@@ D.3.1 Image Understanding
For the initial pre-training stage, we utilized a massive dataset comprising approximately 22M images paired with captions. These data were sourced from a diverse collection of large-scale datasets, including ImageNet (Russakovsky et al., 2014), DataComp-1B (Gadre et al., 2023), COYO-700M (Byeon et al., 2022), SA-1B (Kirillov et al., 2023), and DenseFusion (Li et al., 2024c).
@@
In Stage II, we continued to utilize the comprehensive collection of data from the sources employed in Stage I (ImageNet (Russakovsky et al., 2014), DataComp-1B (Gadre et al., 2023), COYO-700M (Byeon et al., 2022), SA-1B (Kirillov et al., 2023), and DenseFusion (Li et al., 2024c)) for further training.
@@
For the final high-quality fine-tuning stage, we directly utilized the refined 3.2M instruction-tuning data from LLaVA-One-Vision (Li et al., 2024a) to maximize image understanding performance.
@@ D.3.2 Image Generation
For the initial generation training stage, we curated a large-scale dataset totaling approximately 78M samples. This comprises SA-1B (Kirillov et al., 2023), CC-12M (Changpinyo et al., 2021), 1M images from DenseFusion (Li et al., 2024c), and a internal dataset consisting of 55M real images.
@@
Stage II expands upon the data used in Stage I. In addition to SA-1B (Kirillov et al., 2023), CC-12M (Changpinyo et al., 2021), the 55M internal real images, and DenseFusion-1M (Li et al., 2024c), we incorporated 10M synthetic images generated through Flux distillation, following the methodology described in (Shen et al., 2025).
@@
For high-quality generation fine-tuning, we utilized the JourneyDB dataset (Sun et al., 2023) and DenseFusion-1M (Li et al., 2024c). Additionally, we refined the 10M Flux-distilled data from Stage II through filtering, selecting a high-quality subset of 5M images for this final stage.
@@ Appendix E Additional main results
Table 7: Detailed image generation results on the GenEval benchmark. refers to fintuning with GPT-4o distilled dataset (Chen et al., 2025a).
@@
Models Size Single Object Two Objects Count Colors Position Color Attribute Overall Generation-only Models SD3-Med (Esser et al., 2024) 2B 0.99 0.94 0.72 0.89 0.33 0.60 0.74 FLUX.1 [Dev] (Labs et al., 2025) 12B 0.98 0.93 0.75 0.93 0.68 0.65 0.82 DALL-E 3 (Ramesh et al., 2021) - 0.96 0.87 0.47 0.83 0.43 0.45 0.67 Composite Unified Multimodal Models TokenFlow-XL (Qu et al., 2025) 14B 0.95 0.60 0.41 0.81 0.16 0.24 0.55 SEED-X (Ge et al., 2024) 17B 0.97 0.58 0.26 0.80 0.19 0.14 0.49 MetaQuery-XL (Pan et al., 2025) 7B - - - - - - 0.80 Blip3-o (Chen et al., 2025a) 8B - - - - - - 0.84 Smaller Native Unified Multimodal Models Show-o (Xie et al., 2024) 1.3B 0.95 0.52 0.49 0.82 0.11 0.28 0.53 Harmon (Wu et al., 2025d) 1.5B 0.99 0.86 0.66 0.85 0.74 0.48 0.76 Show-o2 (Xie et al., 2025a) 1.5B 0.99 0.86 0.55 0.86 0.46 0.63 0.73 HYDRA 1.5B 1.00 0.95 0.75 0.91 0.76 0.77 0.86 Larger Native Unified Multimodal Models Ming-UniVision (Huang et al., 2025) 16B 1.00 0.93 0.59 0.93 0.92 0.70 0.85 BAGEL (Deng et al., 2025b) 14B 0.98 0.95 0.84 0.95 0.78 0.77 0.88 MUSE-VL (Xie et al., 2025b) 7B 0.98 0.64 0.54 0.72 0.25 0.31 0.57 Janus-Pro (Chen et al., 2025b) 7B 0.99 0.89 0.59 0.90 0.79 0.66 0.80 Show-o2 (Xie et al., 2025a) 7B 1.00 0.87 0.58 0.92 0.52 0.62 0.76 HYDRA 7B 1.00 0.97 0.68 0.91 0.81 0.80 0.86
Table 8: Detailed image generation results on the DPG-Bench benchmark (Hu et al., 2024).
@@
Models Size Global Entity Attribute Relation Other Overall Generation-only Models FLUX.1 [Dev] (Labs et al., 2025) 12B 82.10 89.50 88.70 91.10 89.40 84.00 Smaller Native Unified Multimodal Models Show-o2 (Xie et al., 2025a) 1.5B 87.53 90.38 91.34 90.30 91.21 85.02 HYDRA 1.5B 89.65 91.62 90.24 91.59 90.88 85.51 Larger Native Unified Multimodal Models Janus-Pro (Chen et al., 2025b) 7B 86.90 88.90 89.40 89.32 89.48 84.19 BAGEL (Deng et al., 2025b) 14B 88.94 90.37 91.29 90.82 88.67 85.07 Show-o2 (Xie et al., 2025a) 7B 89.00 91.78 89.96 91.81 91.64 86.14 HYDRA 7B 90.22 91.88 89.92 91.83 90.93 86.40
Table 9: Detailed image generation results on the WISE benchmark (Niu et al., 2025).
@@
Models Size Culture Time Space Biology Physics Chemistry Overall Generation-only Models FLUX.1 [Dev] (Labs et al., 2025) 12B 0.48 0.58 0.62 0.42 0.51 0.35 0.50 SD3.5-Large (Esser et al., 2024) 8B 0.44 0.50 0.58 0.44 0.52 0.31 0.46 Composite Unified Multimodal Models Blip3-o-4B (Chen et al., 2025a) 4B - - - - - - 0.50 Smaller Native Unified Multimodal Models Harmon (Wu et al., 2025d) 1.5B 0.38 0.48 0.52 0.37 0.44 0.29 0.41 Show-o2 (Xie et al., 2025a) 1.5B 0.33 0.41 0.53 0.35 0.49 0.31 0.39 HYDRA 1.5B 0.49 0.47 0.64 0.42 0.59 0.34 0.49 Larger Native Unified Multimodal Models BAGEL (Deng et al., 2025b) 14B 0.44 0.55 0.68 0.44 0.60 0.39 0.52 VILA-U-7B (Wu et al., 2024) 7B 0.26 0.33 0.37 0.35 0.39 0.23 0.31 Janus-Pro-7B (Chen et al., 2025b) 7B 0.30 0.37 0.49 0.36 0.42 0.26 0.35 Emu3-Gen-8B (Wang et al., 2024) 8B 0.34 0.45 0.48 0.41 0.45 0.27 0.39 Show-o2 (Xie et al., 2025a) 7B 0.40 0.45 0.58 0.39 0.53 0.34 0.44 HYDRA 7B 0.52 0.53 0.68 0.47 0.58 0.38 0.53